OpenStudy (anonymous):

what's the definite integral 1/x from [-1,1] zero?

OpenStudy (anonymous):

well it should be zero but both the ∫−101/xdx and the ∫011/xdx are improper integral which are both of divergence?

OpenStudy (anonymous):

well it should be zero but both the \[\int\limits_{-1}^{0}1/xdx\] and the \[\int\limits_{0}^{1}1/xdx\] are improper integral which are both of divergence

OpenStudy (anonymous):

it must be zero right? cause it's symmetric..

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}1/x\]-\[\lim_{y \rightarrow 0}1/y\] =0? same question

OpenStudy (anonymous):

I guess it's inf

OpenStudy (anonymous):

but it's totally symmetric.. you rotate the graph it will offset itself perfectly right?

OpenStudy (anonymous):

ln (1) - ln(-1) how to compute it ?

OpenStudy (anonymous):

 THE definite integral 1/x from [-1,1] means area betweeny=1/x and x axis from x=-1to x=1 from figure it is clear the curve of y=1/x is symmetrical hence both area is equal and has opposite sign if i find total area i.e A1+A2 then we find that its value is zero hence the definite integeral must be zero.....